The perimeter of a rectangle is 68 cm. If the width is increased by 2cm. and twice th length is increased by 10cm, the new perimeter is 100 cm. What are the dimensions of the original rectangle?
Answer by mananth(16949)

let width be x
length =y
2(x+y)=68
x+y=34-------------------1
..
2*((x+2)+(2y+10)=100
x+2+2y+10=50
x+2y+12=50
x+2y=50-12
x+2y=38-----------------2
Subtract equation 1 from 2
x+2y-x-y=38-34
y=4..
x+y=34
plug the value of y
x+4=34
x=30
dimensions are 30 cm by 4 cm